In a group of 40 subjects screened for non￾communicable diseases, 10 are healthy and every 

person of the remaining 30 has either high blood 

pressure, a high level of cholesterol or both. You 

were further informed that 15 have high blood 

pressure and 25 have a high level of cholesterol. 

If a person is selected randomly from this group of 

subjects, what is the probability that he or she has 

either high blood pressure or high level of 

cholesterol?


1
Expert's answer
2021-03-31T13:48:26-0400

25 of the 30 subjects have a high level of cholesterol, so the remaining 5 have high blood pressure, but not a high level of cholesterol. A total of 15 subjects have high blood pressure, 5 of them do not have a high level of cholesterol, hence, the remaining 10 subjects have both high blood pressure and a high cholesterol. Therefore, 25-10=15 subjects have high cholesterol, but not high blood pressure.

At total 5 +15=20 subjects have either high blood pressure or a high level of cholesterol.

If a person is selected randomly from this group of subjects, then the probability that he or she has either high blood pressure or high level of cholesterol is 20/40=0.5.